OpenStudy (anonymous):

What is the value of z in the equation z + 7.8 = 16.4?

OpenStudy (anonymous):

this is a regular equation with the variable of z correct?

OpenStudy (anonymous):

8.6

OpenStudy (anonymous):

Solve for z. What will be the first step into solving this?

OpenStudy (anonymous):

z + 7.8 = 16.4 z = 16.4 - 7.8 z = 8.6